# Did I Faithfully Say What I Thought? Bridging the Gap Between LLMs Neural Activity and Self-Explanations

### v3

@@ Abstract
Large Language Models (LLMs) can generate plausible free text self-explanations to justify their answers. However, these natural language explanations may not accurately reflect the model’s actual reasoning process, indicating a lack of faithfulness. Existing faithfulness evaluation methods rely primarily on behavioral tests or computational block analysis without examining the semantic content of internal neural representations. This paper proposes NeuroFaith, a flexible framework that measures the faithfulness of LLM free text self-explanation by identifying key concepts within explanations and mechanistically testing whether these concepts actually influence the model’s predictions. We show the versatility of NeuroFaith across 2-hop reasoning and classification tasks. Additionally, a linear faithfulness probe based on NeuroFaith is developed to detect unfaithful self-explanations from representation space and improve faithfulness through steering. NeuroFaith provides a principled approach to evaluating and enhancing the faithfulness of LLM free text self-explanations, addressing critical needs for trustworthy AI systems.
@@ 1 Introduction
Autoregressive Large Language Models (LLMs) can generate plausible self Natural Language Explanation (self-NLE) to support their answers (Wiegreffe & Marasovic, ; Huang et al., 2023). Generating self-NLE consists of prompting the LLM to generate an explanation in a predict-then-explain setting, where the model first generates a response to a question and then produces a self-NLE as a justification. Unlike their non-generative predecessors, modern LLMs are trained to generate both answers and free text self-NLE that appear credible despite potentially containing persuasive hallucinations (Sahoo et al., 2024). This way, despite their logical and coherent appearance favoring trust in the model (Han et al., 2023), LLM-generated self-NLE turn out to not systematically reflect the actual underlying decision-making process of the model, creating a tension between self-NLE plausibility and faithfulness (Agarwal et al., 2024).
@@
Faithfulness, as defined by Jacovi & Goldberg (2020), measures ”how accurately the explanation reflects the true reasoning process of the model”, a definition widely adopted in the literature (Lyu et al., 2024) and which we likewise follow throughout this work. Unfaithful self-NLE can have serious consequences in critical domains, where explanations that appear plausible but lack faithfulness might lead end-users to over-rely on model predictions and make unfair (Luo et al., 2022) or harmful (Kayser et al., 2024) decisions. As the use of LLMs is expanding across diverse fields, the combination of their widespread adoption and the simplicity of generating self-NLE through prompting make evaluating their faithfulness increasingly critical.
@@
Assessing self-NLE faithfulness presents profound difficulties. Unlike more structured explainability methods such as attribution (Wiegreffe et al., 2021) or counterfactual approaches (Madsen et al., 2024), the free-form nature of natural language explanations makes faithfulness evaluation particularly difficult. Numerous methods have been developed to measure self-explanation faithfulness (see e.g. Lyu et al. (2024)). However, (1) they mostly perform behavioral tests and do not examine LLM internal reasoning processes (Atanasova et al., 2023; Siegel et al., 2024, 2025; Matton et al., 2025), and (2) they identify the computational blocks that contribute to prediction and self-NLE without conducting semantic analysis of the neural representations within these blocks (Wiegreffe et al., 2021; Parcalabescu & Frank, 2024; Yeo et al., 2024). These shortcomings have led these methods to be characterized as measuring self-consistency rather than genuine faithfulness (Parcalabescu & Frank, 2024), since they fail to establish direct connections between explanations and the model’s actual reasoning processes. To overcome these limitations, we introduce NeuroFaith, a flexible framework for directly measuring LLM self-explanation faithfulness. Our main contributions are as follows:
@@
    We demonstrate the versatility of NeuroFaith by applying it to both 2-hop reasoning tasks and classification problems.
@@
This paper is organized as follows: Section 2 presents how existing approaches measure the faithfulness of LLM self-NLE. Section 3 introduce the NeuroFaith framework. In Sections 4 and 5, we instantiate NeuroFaith for two different tasks, 2-hop reasoning and classification respectively. We finally show in Section 6 that self-NLE faithfulness, as measured by NeuroFaith, can be linearly detected in LLM representation space and improved through steering.
@@ 2 Related Work
Another way to measure self-NLE faithfulness is to compute post-hoc Attribution Agreement (AA) between the prediction and the NLE (Parcalabescu & Frank, 2024; Wiegreffe et al., 2021; Yeo et al., 2024). AA methods compute attribution scores for both the model’s predictions and its self-NLE and then measure the correlation between these scores to assess faithfulness. Higher correlation values indicate greater faithfulness in the model’s self-NLE. AA approaches vary in the post-hoc attribution method employed (gradient-based (Sundararajan et al., 2017), SHAP (Lundberg & Lee, 2017) or activation patching (Meng et al., 2022)). While AA methods assess whether the same LLM computational blocks were used when generating both the prediction and the self-NLE, they overlook the semantic content of neural representations, leaving the model’s reasoning process only partially treated.
@@ 3 NeuroFaith: A Framework for Measuring the Faithfulness of LLM self-NLE
This section introduces the core principles of NeuroFaith, our proposed flexible framework for measuring the faithfulness of LLM self-natural language explanations (self-NLE). Based on the premise that faithful explanations should accurately reflect the model’s internal reasoning process, NeuroFaith quantifies how well a self-NLE aligns with the model’s internal reasoning processes by extracting concepts from self-NLE and mechanistically evaluating their importance for the model prediction. Sections 4 and 5 provide detailed instantiations of how to apply NeuroFaith to 2-hop reasoning and classification tasks respectively. In the following, we denote an $L$-layer auto-regressive Transformer-based LLM $f$, a set of input texts X and a text of interest $x\in\textbf{X}$. We denote $f(x)$ as the model’s answer and $e(x)$ as the corresponding self-NLE produced by $f$ for input $x$. NeuroFaith is a 3-step framework illustrated in Figure 1, summarized below and detailed in the next subsections.
@@
    Concept Extraction. We extract from $e(x)$ a set of concepts $\left\{c_{i}\right\}_{i=1}^{p}$ that influence $f(x)$.
@@
    Faithfulness Measurement. We compute faithfulness $F(x,e)$ by validating that the concepts $\left\{c_{i}\right\}_{i=1}^{p}$ emphasized in the self-NLE actually have significant mechanistic effects on $f(x)$ as determined by their interpretations $\left\{\texttt{I}_{\Gamma}(c_{i})\right\}_{i=1}^{p}$.
@@ 3.1 Concept Extraction
Concept Definition. Recent research has shown that interpretable binary high-level features, referred to as concepts, appear to be linearly encoded within LLM representation space (Elhage et al., 2022; Park et al., 2024). This computational understanding aligns with definitions from cognitive science (Ruiz Luyten & van der Schaar, 2024), where concepts are considered as mental entities essential to thought, enabling information integration and categorization (Goguen, 2005). This alignment makes concepts an ideal granularity level for model interpretation. For example in Figure 1, the concept ”Ingmar Bergman” directly influences the prediction ”Sweden” because Ingmar Bergman is Swedish and directed the movie Persona.
Concept Extraction. Concepts can be extracted either through human investigation or using an auxiliary LLM under LLM-as-a-Judge settings (Gu et al., 2024). Human annotation enables targeted analysis of specific concepts that are hypothesized to be important for the model’s reasoning. Human annotation provides high-quality, domain-expert identification of relevant influential concepts but is costly and may not scale to large datasets. LLM-as-a-Judge approaches offer scalability but may introduce systematic biases or miss relevant concepts that human experts would identify. We provide detailed examples of prompts used to extract concepts from self-NLE with an auxiliary LLM in Appendix D.1.
Figure 1: NeuroFaith overview. NeuroFaith (1) extracting concepts from the self-NLE and (2) assessing the mechanistic influence of these concepts to finally (3) measure faithfulness.
@@ 3.2 Concept-wise Mechanistic Interpretation
Rather than examining every network component, we focus our mechanistic analysis on computational subgraphs that most significantly influence $f$ prediction named circuits (Elhage et al., 2021). Circuits are sparse oriented subgraphs with interpretable functional roles, where nodes represent computational units and edges represent computation paths (Räuker et al., 2023). Formally, we define a circuit as $\Gamma=(\left\{(k,\ell)\right\},\mathcal{G})$ where $\left\{(k,\ell)\right\}$ specifies coordinate pairs (token index, layer) for information flow in $f$, and $\mathcal{G}$ defines the granularity of each computational node. The granularity $\mathcal{G}$ represents transformer sublayers: residual stream output (RS), multi-head attention (MHA), or multi-layer perceptron (MLP(Rai et al., 2024). This granularity avoids focusing on individual neurons due to their polysemanticity (Elhage et al., 2022). For example in Figure 1, the circuit is highlighted in blue, starting at token index 12 and layer 1 and finishes at token index 13 and layer L-1. Circuits can be obtained through manual investigation of task-specific patterns (Bereska & Gavves, 2024; Wang et al., 2023; Biran et al., 2024) or automated discovery via activation patching (Meng et al., 2022; Conmy et al., 2023), backward attribution (Ferrando & Voita, 2024), or Transcoders (Dunefsky et al., 2024). Additional details about circutis are in Appendix D.2.
@@
To mechanistically interpret a concept $c$ across circuit $\Gamma$, we use an interpreter I to compute an attribution score $\texttt{I}_{\Gamma}(c)$ that evaluates whether $c$ impacts $f(x)$. NeuroFaith implements two approaches, each suited to different objectives. Probing methods determine whether $c$ is represented within $f$ hidden states on $\Gamma$. This involves either generating natural language interpretations of hidden states (using Selfie (Chen et al., 2024) or Patchscopes (Ghandeharioun et al., 2024)) to detect $c$, or training linear probes (Belinkov, 2022) when $c$ is linearly separable with sufficient labels to train the probe. Denoting $h^{\ell}_{k}$ the hidden state at token $k$ and layer $\ell$, we define $\texttt{I}_{\Gamma}(c)=\max_{(k,\ell)\in\Gamma}\texttt{I}(h^{\ell}_{k})$, assessing $c$ as important if detected in at least one hidden state from $\Gamma$. Concept importance methods approximate the causal influence of $c$ on $f(x)$ (Geiger et al., 2025). They use gradient-based methods like TCAV (Kim et al., 2018) or Representation Engineering (Zou et al., 2023) to directly manipulate concept-related activations across $\Gamma$ and measure behavioral changes, providing causal evidence for concept importance. Section 5 details this procedure for measuring faithfulness in the case of classification.
@@ 3.3 Faithfulness Measurement
We consider the self-NLE $e(x)$ as faithful when the extracted concepts $\left\{c_{i}\right\}_{i=1}^{p}$ are demonstrably important for $f$’s internal processing according to $\texttt{I}_{\Gamma}(c)$. We propose to define the faithfulness of $e(x)$ as the proportion of concepts it contains that are mechanistically assessed as important, i.e. that have a positive attribution score: $F(x,e)=\frac{1}{p}\sum_{i=1}^{p}\mathbf{1}_{\texttt{I}_{\Gamma}(c_{i})>0}$.
The faithfulness score has different meanings depending on the chosen interpreter I: Probing-based faithfulness measures the detection rate of explanation concepts within the model’s internal representations along circuit $\Gamma$. High probing-based faithfulness indicate that most mentioned concepts are properly decoded from the model’s hidden states. Importance-based faithfulness estimate causal relevance of explanation concepts for the prediction $f(x)$. High importance-based faithfulness indicate that most mentioned concepts actually influence $f$ reasoning process.This framework directly captures concept-level alignment between the self-NLE $e(x)$ and $f$ internal reasoning process. NeuroFaith faithfulness scores either indicate that the self-NLE accurately reflects internal processing or that the explanation mentions concepts that are not mechanistically important. This way, NeuroFaith directly aligns with the established definition of explanation faithfulness introduced above (Jacovi & Goldberg, 2020), offering a principled evaluation of self-NLE faithfulness.
@@ 4 The Case of 2-Hop Reasoning
In the previous section, we defined the high level core principle of NeuroFaith. We now instantiate NeuroFaith to 2-hop reasoning using probing-based faithfulness, and precisely characterize self-NLE with respect to both its faithfulness and its correctness.

### v4

# NeuroFaith: Evaluating LLM Self-Explanation Faithfulness via Internal Representation Alignment

@@ Abstract
Large Language Models (LLMs) can generate plausible free text self-explanations to justify their answers. However, these natural language explanations may not accurately reflect the model’s actual reasoning process, pinpointing a lack of faithfulness. Existing faithfulness evaluation methods rely primarily on behavioral tests or computational block analysis without examining the semantic content of internal neural representations. This paper proposes NeuroFaith, a flexible framework that measures the faithfulness of LLM free text self-explanation by identifying key concepts within explanations and mechanistically testing whether these concepts actually influence the model’s predictions. We show the versatility of NeuroFaith across 2-hop reasoning and classification tasks. Additionally, we develop a linear faithfulness probe based on NeuroFaith to detect unfaithful self-explanations from representation space and improve faithfulness through steering. NeuroFaith provides a principled approach to evaluating and enhancing the faithfulness of LLM free text self-explanations, addressing critical needs for trustworthy AI systems.
Machine Learning, ICML
@@ 1 Introduction
Autoregressive Large Language Models (LLMs) can generate plausible self Natural Language Explanation (self-NLE) to support their answers (Wiegreffe and Marasovic, 2021; Huang et al., 2023). Generating self-NLE consists of prompting the LLM to output an explanation in a predict-then-explain setting, where the model first generates a response to a question and then produces a self-NLE as a justification. Unlike their non-generative predecessors, modern LLMs are trained to generate both answers and free text self-NLE that appear credible despite potentially containing persuasive hallucinations (Sahoo et al., 2024). This way, despite their logical and coherent appearance favoring trust in the model (Han et al., 2023), LLM-generated self-NLE turn out to not systematically reflect the actual underlying decision-making process of the model, creating a tension between self-NLE plausibility and faithfulness (Agarwal et al., 2024).
Figure 1: 2-hop reasoning example with LLM answer and decoded hidden representation ($\tilde{h}$). Faithfulness linear probe is applied before ($e$) and after ($e_{steer}$) steering. Red: unfaithful activations (e.g. Morris Halle); Green: faithful activations (e.g. Noam Chomsky).
@@
Faithfulness, as defined by Jacovi and Goldberg (2020), measures ”how accurately the explanation reflects the true reasoning process of the model”, a definition widely adopted in the literature (Lyu et al., 2024) and which we also follow throughout this work. Unfaithful self-NLE can have serious consequences in critical domains (Farah et al., 2023), where explanations that appear plausible but lack faithfulness might lead end-users to over-rely on model predictions and make unfair (Luo et al., 2022) or harmful (Kayser et al., 2024) decisions. As the use of LLMs is expanding across diverse fields, the combination of their widespread adoption and the simplicity of generating self-NLE through prompting make evaluating their faithfulness increasingly critical.
@@
Assessing self-NLE faithfulness presents profound difficulties due to the free-form nature of natural language explanations, unlike more structured explainability methods such as attribution (Wiegreffe et al., 2021) or counterfactual approaches (Madsen et al., 2024). Numerous methods have been developed to measure self-explanation faithfulness (see e.g. Lyu et al. (2024)). However, (1) they mostly perform behavioral tests and do not examine LLM internal reasoning processes (Atanasova et al., 2023; Siegel et al., 2024, 2025; Matton et al., 2025), and (2) they identify the computational blocks that contribute to prediction and self-NLE without conducting semantic analysis of the neural representations within these blocks (Wiegreffe et al., 2021; Parcalabescu and Frank, 2024; Yeo et al., 2025). These shortcomings have led these methods to be characterized as measuring self-consistency rather than genuine faithfulness (Parcalabescu and Frank, 2024), since they fail to establish direct connections between explanations and the model’s reasoning processes. To overcome these limitations, we introduce NeuroFaith, a flexible framework for directly measuring LLM self-NLE faithfulness. Our main contributions are as follows:
@@
    We demonstrate the versatility of NeuroFaith by applying it to both 2-hop reasoning and classification tasks.
@@
”Figure 1 illustrates NeuroFaith on a 2-hop reasoning question: despite answering correctly, the model’s initial self-NLE ($e$) is unfaithful. Linear steering improves faithfulness by aligning the steered self-NLE ($e_{steer}$) with the decoded internal concept (Noam Chomsky) being required for reasoning solving. This paper is organized as follows: Section 2 presents how existing approaches measure the faithfulness of LLM self-NLE. Section 3 introduce the NeuroFaith framework. In Sections 4 and 5, we instantiate NeuroFaith for two different tasks, 2-hop reasoning and classification respectively. We finally show in Section 6 that self-NLE faithfulness, as measured by NeuroFaith, can be linearly detected in LLM representation space and improved through steering.
@@ 2 Related Work
Another way to measure self-NLE faithfulness is to compute post-hoc Attribution Agreement (AA) between the prediction and the NLE (Parcalabescu and Frank, 2024; Wiegreffe et al., 2021; Yeo et al., 2025). AA methods compute attribution scores for both the model’s predictions and its self-NLE and then measure the correlation between these scores to assess faithfulness. Higher correlation values indicate greater faithfulness in the model’s self-NLE. AA approaches vary in the post-hoc attribution method employed (gradient-based (Sundararajan et al., 2017), SHAP (Lundberg and Lee, 2017) or activation patching (Meng et al., 2022)). While AA methods assess whether the same LLM computational blocks were used when generating both the prediction and the self-NLE, they overlook the semantic content of neural representations, leaving the model’s reasoning process only partially treated.
@@ 3 NeuroFaith: A Framework for Measuring the Faithfulness of LLM self-NLE
This section introduces the core principles of NeuroFaith, our proposed flexible framework for measuring the faithfulness of LLM self-natural language explanations (self-NLE). Based on the premise that faithful explanations should accurately reflect the model’s internal reasoning process, NeuroFaith quantifies how well a self-NLE aligns with the model’s internal reasoning processes by extracting concepts from self-NLE and mechanistically evaluating their importance for the model prediction. Sections 4 and 5 provide detailed instantiations of how to apply NeuroFaith to 2-hop reasoning and classification tasks respectively. In the following, we denote an $L$-layer auto-regressive Transformer-based LLM $f$, a set of input texts X and a text of interest $x\in\textbf{X}$. We denote $f(x)$ the model’s answer and $e(x)$ the corresponding self-NLE produced by $f$ for input $x$. NeuroFaith is a 3-step framework illustrated in Figure 2, summarized below and detailed in the next subsections.
@@
    Concept Extraction. We extract from $e(x)$ a set of concepts $\left\{c_{i}\right\}_{i=1}^{p}$ quoted as important for $f(x)$.
@@
    Faithfulness Measurement. We compute faithfulness $F(x,e)$ by validating that the detected concepts $\left\{c_{i}\right\}_{i=1}^{p}$ in the self-NLE have mechanistic effects on $f(x)$ as determined by their interpretations $\left\{\texttt{I}_{\Gamma}(c_{i})\right\}_{i=1}^{p}$.
@@ 3.1 Concept Extraction
Concept Definition. Recent research has shown that interpretable binary high-level features, referred to as concepts, appear to be linearly encoded within LLM representation space (Elhage et al., 2022; Park et al., 2024). This computational understanding aligns with definitions from cognitive science (Ruiz Luyten and van der Schaar, 2024), where concepts are considered as mental entities essential to thought, enabling information integration and categorization (Goguen, 2005). This alignment makes concepts an ideal granularity level for model interpretation. For example in Figure 2, the concept ”Ingmar Bergman” directly influences the prediction ”Sweden” because Ingmar Bergman is Swedish and directed the movie Persona.
Concept Extraction. Concepts can be extracted either through human investigation or using an auxiliary LLM under LLM-as-a-Judge settings (Gu et al., 2024). Human annotation enables targeted analysis of specific concepts that are hypothesized to be important for the model’s reasoning. Human annotation provides high-quality, domain-expert identification of relevant influential concepts but is costly and may not scale to large datasets. LLM-as-a-Judge approaches offer scalability but may miss relevant concepts that human experts would identify. We provide detailed examples of prompts used to extract concepts from self-NLE with an auxiliary LLM in Appendix D.1.
Figure 2: NeuroFaith overview. NeuroFaith (1) extracts concepts from the self-NLE (Section 3.1), (2) assesses the mechanistic influence of these concepts (Section 3.2) to finally (3) measure faithfulness (Section 3.3).
@@ 3.2 Concept-wise Mechanistic Interpretation
Rather than examining every network component, we focus our mechanistic analysis on computational subgraphs that most significantly influence $f$ prediction named circuits (Elhage et al., 2021). Circuits are sparse oriented subgraphs with interpretable functional roles, where nodes represent computational units and edges represent computation paths (Räuker et al., 2023). Formally, we define a circuit as $\Gamma=(\left\{(k,\ell)\right\},\mathcal{G})$ where $\left\{(k,\ell)\right\}$ specifies coordinate pairs (token index, layer) for information flow in $f$, and $\mathcal{G}$ defines the granularity of each computational node. The granularity $\mathcal{G}$ represents transformer sublayers: residual stream output (RS), multi-head attention (MHA), or multi-layer perceptron (MLP(Rai et al., 2024). This granularity avoids focusing on individual neurons due to their polysemanticity (Elhage et al., 2022). For example in Figure 2, the circuit is highlighted in blue, starting at token index 12 and layer 1 and finishes at token index 13 and layer L. Circuits can be obtained through manual investigation of task-specific patterns (Bereska and Gavves, 2024; Wang et al., 2023; Biran et al., 2024) or automated discovery via activation patching (Meng et al., 2022; Conmy et al., 2023), backward attribution (Ferrando and Voita, 2024), or Transcoders (Dunefsky et al., 2024). Additional details about circutis are in Appendix D.2.
@@
To mechanistically interpret a concept $c$ and its impact on $f(x)$ across circuit $\Gamma$, we use an interpreter I to compute an attribution score $\texttt{I}_{\Gamma}(c)$ that evaluates whether $c$ impacts $f(x)$. NeuroFaith implements two approaches, each suited to different objectives. Probing methods determine whether $c$ is represented within $f$ hidden states in $\Gamma$. This involves either generating natural language interpretations of hidden states (using Selfie (Chen et al., 2024) or Patchscopes (Ghandeharioun et al., 2024)) to detect $c$, or training linear probes (Belinkov, 2022) when $c$ is linearly separable with sufficient labels to train the probe. Denoting $h^{\ell}_{k}$ the hidden state at token $k$ and layer $\ell$, we define $\texttt{I}_{\Gamma}(c)=\max_{(k,\ell)\in\Gamma}\texttt{I}(h^{\ell}_{k})$, assessing $c$ as important if detected in at least one hidden state from $\Gamma$. Concept importance methods approximate the causal influence of $c$ on $f(x)$ (Geiger et al., 2025). They use gradient-based methods like TCAV (Kim et al., 2018) or Representation Engineering (Zou et al., 2023) to directly manipulate concept-related activations across $\Gamma$ and measure behavioral changes, providing causal evidence for concept importance. Section 5 details this procedure in the case of classification.
@@ 3.3 Faithfulness Measurement
We consider the self-NLE $e(x)$ as faithful when the extracted concepts $\left\{c_{i}\right\}_{i=1}^{p}$ are demonstrably important for $f$’s internal processing according to $\texttt{I}_{\Gamma}(c)$. We propose to define the faithfulness of $e(x)$ as the proportion of concepts it contains being mechanistically important, i.e. having positive attribution score: $F(x,e)=\frac{1}{p}\sum_{i=1}^{p}\mathbf{1}_{\texttt{I}_{\Gamma}(c_{i})>0}$.
The faithfulness score has different meanings depending on the chosen interpreter I: Probing-based faithfulness measures the detection rate of explanation concepts within the model’s internal representations along circuit $\Gamma$. High probing-based faithfulness indicate that most mentioned concepts are properly decoded from the model’s hidden states. Importance-based faithfulness estimate causal relevance of explanation concepts for the prediction $f(x)$. High importance-based faithfulness indicate that most mentioned concepts actually influence $f$ reasoning process. This framework directly captures concept-level alignment between the self-NLE $e(x)$ and $f$ internal reasoning process. NeuroFaith faithfulness scores either indicate that the self-NLE accurately reflects internal processing or that the explanation mentions concepts that are not mechanistically important. This way, NeuroFaith directly aligns with the established definition of explanation faithfulness introduced above (Jacovi and Goldberg, 2020), offering a principled evaluation of self-NLE faithfulness.
Table 1: 2-hop reasoning accuracy, self-NLE correctness, latent first hop correctness and self-NLE faithfulness across models obtained from NeuroFaith. on the the Wikidata-2-hop dataset. ”(In)accurate” represents (in)correct predictions.
Model Task Accuracy Self-NLE Latent Hop 1 Self-NLE
Correctness Correctness Faithfulness
Accurate Inaccurate Accurate Inaccurate Accurate Inaccurate
gemma-2-2b 7.1% 58.0% 56.0% 48.3% 47.4% 48.4% 57.6%
gemma-2-9b 16.6% 74.0% 55.4% 58.5% 44.3% 60.8% 55.0%
gemma-2-27b 22.3% 77.8% 55.6% 64.9% 46.9% 68.8% 60.2%
mistral-3-7b 21.2% 70.2% 37.9% 52.9% 29.6% 51.1% 32.6%
@@ 4 The Case of 2-Hop Reasoning
In the previous section, we defined the high level core principle of NeuroFaith. We now instantiate NeuroFaith to 2-hop reasoning using probing-based faithfulness.